OpenStudy (anonymous):

(will give medal) acceleration is defined as the rate of change for which of the following? A. time B. velocity C. position Displacement

OpenStudy (solomonzelman):

acceleration is the slope (the derivative) of the velocity

OpenStudy (solomonzelman):

and slope is the same thing as rate of change --- just in case you didn't know.

OpenStudy (anonymous):

ok so can you elaborate on that a little more plz?

OpenStudy (solomonzelman):

lets define: displacement as \(\large\color{black}{ \displaystyle s(t) }\) velocity as \(\large\color{black}{ \displaystyle s'(t) }\) \(\large\color{black}{ \displaystyle (}\) or v(t) \(\large\color{black}{ \displaystyle )}\) acceleration as \(\large\color{black}{ \displaystyle s''(t) }\) \(\large\color{black}{ \displaystyle (}\) or a(t) \(\large\color{black}{ \displaystyle )}\) Time is just what the x-scale. ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

OpenStudy (solomonzelman):

velocity is the slope of the displacement acceleration is the slope of velocity acceleration is slope-of-slope of displacement

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle s'(t) }\) is the notation for the 1st derivative, or the first time slope of \(\large\color{black}{ \displaystyle s(t) }\). \(\large\color{black}{ \displaystyle s''(t) }\) is the notation for the 2st derivative, or the second time slope of \(\large\color{black}{ \displaystyle s(t) }\).

OpenStudy (anonymous):

so the answer is velocity?

OpenStudy (solomonzelman):

Yes

OpenStudy (anonymous):

thanks :) giving you a medal

OpenStudy (solomonzelman):

or inversely, \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}a(t)~dt~=v(t)}\)

OpenStudy (solomonzelman):

I don;t need a medal, but if you want you can